# Ultrafast All-Optical Measurement of Squeezed Vacuum in a Lithium Niobate Nanophotonic Circuit

### v1

@@ I Introduction
A crucial part of any quantum information processing system is a measurement device capable of characterizing a component of interest. For SV, homodyne detection is often employed as its phase-sensitive nature can be used to reconstruct the Wigner function, a quasi-probability distribution over two non-commuting quadratures which completely characterizes a quantum state. Homodyne detectors can isolate a single mode by shaping the spectral profile of the pump appropriately [32]. This phase-sensitive and mode-selective behavior makes them well-suited for characterizing a variety of quantum states, including non-classical states. Single-photon and photon-number-resolving (PNR) detectors are also common tools for state characterization, especially in the context of states exhibiting Wigner negativity [33, 34]. Given the non-Gaussian nature of their measurement operator, PNR detectors have been used to generate non-classical states which are characterized through homodyne detection [35, 36, 37, 38]. While homodyne and PNR detection are powerful techniques for probing quantum states of interest, their speed is ultimately constrained by the bandwidth of the electronics used to physically realize these detectors, which are limited to the GHz range.
@@ II.1 Quadrature Measurement
Measuring $N$ photons over $n$ modes can be expressed with the constructed detection operator [54]:
@@
$j=\{k_{1},...k_{n}\}\mid\sum_{i}^{n}k_{i}=N$ (8)
where the operator |$\ket{k_{i}}\bra{k_{i}}$ | measures $k$ photons in mode $i$. Here, we have constrained the values of $k$ used to construct $\hat{N}$ such that the number of photons measured in each mode sums to $N$. The sum over $j$ captures all possible combinations of $k_{i}$ values which sum to $N$ as the photodetector cannot distinguish between photons arriving from different modes. $P(N,\phi)$ is then defined as
$P(N,\phi)=Tr(\hat{N}\rho_{\phi})$ (9)
where $\rho_{\phi}$ represents the SV after it has been amplified along $\phi$.
At the photodetector, we use a combination of a 1700-nm long-pass filter and 1950-nm short-pass filter to suppress contributions from higher order modes. Using the measured $P(N)$ for vacuum amplification (i.e when no signal field is sent into the OPA), we calculate a Schmidt number of 1.35 modes using definition of $g^{(2)}$ from [55]. Contributions from the higher-order modes are quantified and separated by fitting a 2-mode photon number distribution to determine the gain of the fundamental mode as a function of $\phi$. These distributions are then sampled and used as inputs to a maximum-likelihood reconstruction algorithm to recover the density matrix and Wigner function [56]. This two-mode distribution can be calculated from the convolution of two single-mode distributions [57].
@@ IV.1 Shot-noise Calibration
$S_{\pm}^{\phi}[dB]=10\log_{10}\Big{(}\frac{\braket{N(\phi)}}{\braket{N_{vac}}}\Big{)}$ (10)
@@ IV.3 Pulse-to-pulse Measurements
Figure 6: A) Sample photon number distributions measured at different $\phi$. Red points are data, blue curve is a 2-mode fit. B) Photon number variance vs $\phi$ for slow and fast (pulse-to-pulse) measurements, and the 1st mode of a 2-mode fit. C) Recovered squeezed vacuum Wigner function. D) Expected density matrix ($\rho$) based on measured squeezing and anti-squeezing. E) Recovered $\rho$. F) Absolute difference between expected and recovered $\rho$.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. Sample measured histograms at 4 different phases are plotted in Fig. 6A, and each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a 2-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. To calculate the 2-mode distribution, we start with the single-mode distribution given by [42]:
Convolving this distribution with itself yields the 2-mode distribution [57]:
[…]
Intuitively, this convolution can be understood as a sum of all probabilities for which the photon contributions of the first and second modes sum to $N$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the 2-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon-number offset resulting from pump depletion, the numerator and denominator of Eq.10 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The 2-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm which calculates the Wigner functions for squeezing data (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer strong enough to provide a sufficiently large signal-to-noise ratio at the fast photodetector.

### v2

@@ I Introduction
A crucial part of any quantum information processing system is a measurement device capable of characterizing a component of interest. For SV, homodyne detection is often employed as its phase-sensitive nature can be used to reconstruct the Wigner function, a quasi-probability distribution over two non-commuting quadratures which completely characterizes a quantum state. Homodyne detectors can isolate a single mode by shaping the spectral profile of the pump appropriately [32]. This phase-sensitive and mode-selective behavior makes them well-suited for characterizing a variety of quantum states, including non-classical states. Single-photon and photon number-resolving (PNR) detectors are also common tools for state characterization, especially in the context of states exhibiting Wigner negativity [33, 34]. Given the non-Gaussian nature of their measurement operator, PNR detectors have been used to generate non-classical states which are characterized through homodyne detection [35, 36, 37, 38]. While homodyne and PNR detection are powerful techniques for probing quantum states of interest, their speed is ultimately constrained by the bandwidth of the electronics used to physically realize these detectors, which are limited to the GHz range.
@@ II.1 Quadrature Measurement
At the photodetector, we use a combination of a 1700-nm long-pass filter and 1950-nm short-pass filter to suppress contributions from higher order modes. Using the measured $P(N)$ for vacuum amplification (i.e when no signal field is sent into the OPA), we calculate a Schmidt number of 1.35 modes using definition of $g^{(2)}$ from [54]. Contributions from the higher-order modes are quantified and separated by fitting a 2-mode photon number distribution to determine the gain of the fundamental mode as a function of $\phi$. For a single mode, the photon number distribution is
where $\langle N\rangle(r)$ is the squeezing-dependent average photon number. Convolving this distribution with itself yields the 2-mode distribution [55]:
@@
Intuitively, this convolution can be understood as a sum of all probabilities for which the photon contributions of the first and second modes sum to $N$ as the photodetector cannot distinguish between photons arriving from different modes. These fitted distributions are then sampled and used as inputs to a maximum-likelihood reconstruction algorithm to recover the density matrix and Wigner function [56].
@@ IV.1 Shot-noise Calibration
$S_{\pm}^{\phi}[dB]=10\log_{10}\Big{[}\frac{\braket{N(\phi)}}{\braket{N_{vac}}}\Big{]}$ (9)
@@ IV.3 Pulse-to-pulse Measurements
Figure 6: A) Sample photon number distributions from fast measurements at different $\phi$. Red points are data, blue curve is a 2-mode fit. B) Photon number variance vs $\phi$ for slow and fast (pulse-to-pulse) measurements, and the 1st mode of a 2-mode fit. C) Recovered squeezed vacuum Wigner function. D) Expected density matrix ($\rho$) based on measured squeezing and anti-squeezing. E) Recovered $\rho$. F) Absolute difference between expected and recovered $\rho$.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. Sample measured histograms at 4 different phases are plotted in Fig. 6A, and each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a 2-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the 2-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon number offset resulting from pump depletion, the numerator and denominator of Eq.9 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The 2-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm which calculates the Wigner functions for squeezing data (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998 $\pm$ 0.0001. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer strong enough to provide a sufficiently large signal-to-noise ratio at the fast photodetector.